OpenStudy (jonathan5824):

Jim is designing a seesaw for a children's park. The seesaw should make an angle of 45° with the ground, and the maximum height to which it should rise is 1 meter, as shown below: A see-saw is shown with one end on the ground and the other in the air. The see-saw makes an angle of 45 degrees with the ground. The height of the see-saw from the ground, at the other end, is labeled 1 meter. What is the maximum length of the seesaw? 1 meter 1.4 meters 2 meters 0.5 meters

OpenStudy (wolf1728):

The seesaw, the ground and that "height line" form a 45 45 90 triangle. Do you know the proportions of a 45 45 90 triangle. OR you can use the Pythagorean Theorem.

OpenStudy (jonathan5824):

no I am really confused plz explain

OpenStudy (jonathan5824):

would the proportions be 1/2

OpenStudy (wolf1728):

That drawing isn't very well done. That angle is closer to 30 degrees instead of 45. Nonetheless we can get an answer.

OpenStudy (jonathan5824):

ok

OpenStudy (jonathan5824):

how do we do it

OpenStudy (wolf1728):

As I said the drawing is rather poor. Basically, the distance from the bottom of the seesaw on the left to the "1 meter" height line on the right is one meter.

OpenStudy (jonathan5824):

ok

OpenStudy (wolf1728):

To solve for the hypotenuse, we know the "legs " of the triangle are 1 meter each. Using the Pythagorean Theorem, side 1 ^2 + side 2 ^2 = hypotenuse ^2 hypotenuse^2 = 1 +1 hypotenuse^2 = 2 hypotenuse = sq root of 2

OpenStudy (jonathan5824):

1.41421356

OpenStudy (wolf1728):

yes jonathan5824 that is it

OpenStudy (jonathan5824):

that's the answer

OpenStudy (wolf1728):

yes 1.414 meters

OpenStudy (jonathan5824):

yhea I was right can you help me with one more please you are very helpful

OpenStudy (wolf1728):

Okay - one more (and thanks)

OpenStudy (jonathan5824):

Look at the figure below: An image of a right triangle is shown with an angle labeled y. If sin y° = 7 divided by q and tan y° = 7 divided by r ,what is the value of cos y°? cos y°= r divided by q cos y° = 7r cos y° = 7q cos y°= q divided by r

OpenStudy (wolf1728):

Gee, if you don't have a graphic I'll draw one I think I'll make one

OpenStudy (jonathan5824):

its either A or D pretty sure A though

OpenStudy (wolf1728):

Here's the graphic

OpenStudy (wolf1728):

sin y = 7/q tan y = 7/r sin = opposite/ hypotenuse tangent = opposite/adjacent

OpenStudy (jonathan5824):

ok now what?

OpenStudy (wolf1728):

Here's a new graphic

OpenStudy (wolf1728):

We can say the opposite side = 7 you see why?

OpenStudy (jonathan5824):

It tells you that in the question though

OpenStudy (jonathan5824):

do you agree with A?

OpenStudy (wolf1728):

THe question does not say the opposite side is 7 so hypotenuse = q and adjacent side = r

OpenStudy (jonathan5824):

wait so it would be q/r so D correct?

OpenStudy (wolf1728):

cosine = adjacent / hypotenuse cosine = r / q

OpenStudy (wolf1728):

That would be A

OpenStudy (jonathan5824):

yhea I was correct I have some others but you have helped much thank you

OpenStudy (wolf1728):

u r welcome :-)